Derive the expression X-X0=∫vdt and V-V0=∫adt. and explain..
 

Grade:11

1 Answers

Vikas TU
7 years ago
rate of change of velocity is accln. from the definition.
Thus,
a = dv/dt
dv = a dt
integrating both sides we get,
at = v + c
at t=0 we edfine c = u
Therfore,
v =u + at this is frist law of eqn. of motion.
Simlarly, 
v = dx/dt = u + at
dx = udt + atdt
Integerating both sides we get,
x = ut + 0.5at^2